John invested the following amounts in three stocks: Security Investment Beta Stock A $584,657 1.14 Stock B $680,099 2.03 Stock C $208,590 0.84 Calculate the beta portfolio.
Answer #1

Total investment=(584,657+680,099+208590)=$1,473,346

Beta portfolio=Respective beta*Respective investment weight

=(584657/1,473,346*1.14)+(680,099/1,473,346*2.03)+(208590/1,473,346*0.84)

which is equal to

=1.51(Approx).